# ANN vs SNN: A case study for Neural Decoding in Implantable Brain-Machine Interfaces

### v1

@@ Abstract
While it is important to make implantable brain-machine interfaces (iBMI) wireless to increase patient comfort and safety, the trend of increased channel count in recent neural probes poses a challenge due to the concomitant increase in the data rate. Extracting information from raw data at the source by using edge computing is a promising solution to this problem, with integrated intention decoders providing the best compression ratio. In this work, we compare different neural networks (NN) for motor decoding in terms of accuracy and implementation cost. We further show that combining traditional signal processing techniques with machine learning ones deliver surprisingly good performance even with simple NNs. Adding a block Bidirectional Bessel filter provided maximum gains of $\approx 0.05$, $0.04$ and $0.03$ in $R^{2}$ for ANN_3d, SNN_3D and ANN models, while the gains were lower ($\approx 0.02$ or less) for LSTM and SNN_streaming models. Increasing training data helped improve the $R^{2}$ of all models by $0.03-0.04$ indicating they have more capacity for future improvement. In general, LSTM and SNN_streaming models occupy the high and low ends of the pareto curves (for accuracy vs. memory/operations) respectively while SNN_3D and ANN_3D occupy intermediate positions. Our work presents state of the art results for this dataset and paves the way for decoder-integrated-implants of the future.
Index Terms:
implantable-brain machine interface, intention decoder, spiking neural networks, low-power.
@@ I Introduction
Implantable Brain-Machine Interfaces (iBMI) (Fig. 1(a)) are a promising class of assistive technology that enables the reading of a person’s intent to drive an actuator[1]. It holds promise to enable paralyzed patients to perform activities of daily living with partial or total autonomy[1]. While the first applications were in motor prostheses to control a cursor on a computer screen[2], or wheelchairs[3], or robotic arms[4], recent studies have shown remarkable results for speech decoding[5, 6], handwritten text generation[7] and therapies for other mental disorders[8].
@@
The majority of clinical iBMI systems have a wired connection from the implant to the outside world [1]. However, this entails a risk of infection leading to the increasing interest in wireless neural interfaces [9, 1, 10]. Another trend in the field has been the continual increase in the number of electrodes[11] to increase the number of simultaneously recorded neurons (Fig. 1(b)) which can increase the accuracy of decoding user intent and enable dexterous control. Recently developed Neuropixels technology has pushed the number of recorded neurons to $\approx 1000$. However, this makes it a problem for wireless implants due to the conflicting requirements of high data rate and low power consumption[12]. Hence, there are efforts to compress the neural data on the sensor by extracting information from it by edge computing (Fig. 1(c)).
@@
Different degrees of computing can be embedded in the implant ranging from spike detection, sorting to decoding[13]. Ideally, decoding on implant can provide the maximum compression [14] with the added benefit of patient privacy since the data does not need to leave the implant—only motor commands are sent out. Traditional decoders have used methods from statistical signal processing such as Kalman Filters and their variants[15]. With the rapid growth of Artificial Neural Networks (ANN) and variants for many different applications, it is natural to explore the usage of such techniques for motor decoding and several such works have recently been published [14, 16].
@@
To fit on the implant, the decoder has to be extremely energy and area efficient along with being accurate. Brain-inspired SNN are supposed to be more energy-efficient due to their event-driven nature. They are also expected to be better at modeling signals with temporal dynamics due to their inherent “stateful” neurons with memory. However, detailed comparisons between SNN and ANN variants with controlled datasets and benchmarking procedures have been lacking. A recent effort[17] has put together a benchmarking suite to address this gap and one chosen task is that of motor decoding. We use the same dataset for benchmarking and show additional results for more control cases.
@@
Neurobench showed that streaming SNNs provide a good tradeoff in terms of accuracy vs computes while other methods could have similar memory footprint. Further, it was shown that expanding traditional ANN with temporal memory (ANN-flat) at the input could drastically increase accuracy albeit at the cost of increased operations. Hence, we ask the question – will ANN models augmented with memory be better than SNNs in terms of the tradeoff between accuracy and cost (memory, operations/energy)? We make the following novel contributions in this paper:
  • We compare SNNs with ANNs with memory augmentation at hidden state (by using LSTM) and output (by incorporating a traditional Bessel filter from signal processing).
  • We show that combining NNs with traditional signal processing methods such as filtering drastically improves performance at minimial cost of additional operations or memory.
@@
  • We show the effect of testing with better curated data.
The rest of the paper is organized as follows. The following section discusses some of the related works while Section III describes the dataset, models and pre-processing used in this work. Section IV presents the results comparing different models in terms of their performance-cost tradeoff using pareto curves. This is followed by a section V that discusses the main results and provides additional control experiments. Finally, we summarize our findings and conclude in the last section.
Figure 1: (a) Conceptual figure of an iBMI that reads user intent and controls an effector. (b) Trend of exponential increase in the number of simultaneously recorded neurons[18]. (c) Integrating computing in the implant can reduce the wireless datarate enabling scalability of iBMI systems.
II Related Works and Contribution
@@
II-A Traditional Signal Processing Decoders
@@
II-B Machine Learning Decoders
@@
Neuromorphic algorithms have emerged as an energy-efficient decoder and an effective tool for data compression [27]. SNN is a brain-inspired neural network popular in neuromorphic applications due to its low energy. It can achieve nearly the same accuracy as ANN but with less than $10\%$ memory access and computation of ANN [28]. Similarly, it was found that the SNN decoders could use far fewer computes compared to ANN, but with a performance penalty in accuracy, for the motor prediction of primates in the Neurobench benchmark suite[17].
In this work, we show that the combination of traditional signal processing and machine learning algorithms results in the best decoders for iBMI systems.
III Methodology
@@
III-A Dataset
The primate reaching dataset chosen for this paper was gathered and released by [29], with the six files chosen for Neurobench [17] being the files of interest. These six files are recordings of two non-human primates (NHP) (Indy and Loco), where each NHP accounts for three files (more details about this choice in [17]). This dataset contains microelectrode array (MEA) recordings of the NHP’s brain activity while it is moving a cursor to the target location, as seen in Figure 2. The finger velocity is sampled at $f_{GT}=250$ Hz resulting in ground truth labels at a fixed interval of $T_{GT}=4$ ms. The target position changes once the monkey successfully moves the cursor to the intended target. We refer to this action as a reach. The dataset contains a continuous stream of the brain’s activity from one MEA with $N_{ch}=96$ probes (Indy) or two MEAs with $N_{ch}=192$ probes (Loco). In this work, we ignore sorted spikes since it has been shown that spike detection provides sufficient information for decoding[30] and is more stable over time. Hence, the number of probes $N_{ch}$ is the input feature dimension $N_{0}$ for the neural network models (except ANN_3D) that will be discussed in the following subsection.
Training NN models on time series-based data requires the data to be split apart into separate segments. In analogy with keyword spotting[17], each segment of neural data should correspond to separate keywords. By using the target positions in this dataset, we can separate the spike data into segments based on indices in the target position array where there is a change in values, as illustrated in Figure 3. Such consecutive indices forms the beginning and end of a reach, and then we can split the time series into training, validation, and test sets based on the number of reaches. The split ratio used in this paper follows that of Neurobench[17], which is 50% for the training set, and 25% each for validation and test sets. The total number of reaches recorded in each file can be seen in Table I.
@@ III-A Dataset
TABLE I: Number of reaches in each file
@@
indy_20160622_01
970
indy_20160630_01
1023
indy_20170131_02
635
loco_20170131_02
587
loco_20170215_02
409
loco_20170301_05
472
III-B Network Models
@@
To explore the potential of various neural network models as the neural decoder, five different model architectures with and without memory are tested: ANN, ANN_3D, SNN_3D, Streaming SNN and LSTM, which can be seen in Figure 4. These five models use popular NN architectures and have memory at the input layer or hidden layer. Every model except for LSTM has two versions of varying complexity (explained in section IV-A) where complexity refers to the model size indicating the number of neurons. The larger model is henceforth referred to as the base model while the smaller model is dubbed the tiny variant. It was found that networks deeper than 3 layers performed poorly and hence deeper models were excluded from this study.
@@
    The ANN model has an architecture of $N_{ch}-N_{1}-N_{2}-2$, with rectified linear unit (ReLU) as the activation function for the first two-layers as well as batch normalization to improve upon the accuracy obtained by the model. Note that $N_{0}=N_{ch}$ indicates one feature extracted from each probe obtained by summing the neural spikes over a fixed duration of $T_{W}$ as described in Section III-C. Also, $N_{3}=2$ corresponds to predicting the $X$ and $Y$ velocities. A dropout layer with a dropout rate of $0.5$ is also added to the first two layers to help regularize the model. In analogy with the naming convention of ANN_3D introduced next, this model can also be referred to as ANN_2D due to the shape of the input weight tensor.
@@
    The architecture of the ANN_3D or ANN_flat model is $m\times N_{ch}-N_{1}-N_{2}-2$, i.e. it shares an identical architecture with ANN, except at the input layer. This model divides the $T_{W}$ duration of the input bin window into $m$ sub-windows and creates a $m$-dimensional feature from each probe by summing spikes in each sub-window. This mode of input will be further explained in Section III-C. The input will then be flattened across the sub-windows, yielding a final input dimension of $N_{ch}\times m$; hence, the number of weights/synapses in the first layer is $m$ times more than ANN. It is referred to as ANN_flat in [17]; we refer to it as ANN_3D here in keeping with the shape of the input weight tensor, which we feel is more intuitive.
@@
    The SNN_3D shares a similar architecture with ANN ($N_{ch}-N_{1}-N_{2}-2$), with the following differences: 1) Instead of using standard activation like ReLU, the SNN_3D model uses the leaky integrate-and-fire (LIF) neuron after every fully-connected layer, 2) the input is first passed through layer normalization, similar to LSTM due to the recurrent nature of LIF, and 3) at the final layer there is a scaling layer applied to the output LIF neurons. The LIF neurons are governed by the following set of equations:
@@
    where $U[t]$ and $X[t]$ are the membrane potential of the LIF neuron and the input at the t-th time step respectively, $W$ is the synaptic weight of the fully-connected layer, $\beta$ is the decay rate, $S_{out}[t]$ is the output spike, $U_{thr}$ is the membrane potential threshold, $U_{sub}$ is the subtracted value if the reset mechanism is reset-by-subtraction and $\theta$ is the reset mechanism. The LIF neurons for all layers shares the same $U_{thr}$ and $\beta$. The first two layer uses the reset-to-zero mechanism while the last layer does not use any reset to allow the final output neurons to accumulate membrane potential to predict the velocity of the primate’s movement. For every stride of $4$ ms, the membrane voltages are reset and the integration is restarted with fresh input to produce the next output.
    The input for the SNN_3D model is different from ANN and identical to ANN_3D; however, unlike the ANN_3D model, the spike counts in the $m$ sub-windows spanning $T_{W}$ are input to the LIF neuron using only one weight/synapse but over $m$ time steps. Due to the reset of the LIF neurons after every prediction, overlapping bin-windows (for $T_{W}>st$) cause the SNN_3D to process the same input spikes for multiple predictions.
@@
    The SNN_Streaming model also consists of three fully-connected layers ($N_{ch}-N_{1}-N_{2}-2$), with LIF neurons (see Eq. 1) in each layer. Unlike SNN_3D, every LIF layer has its own unique $U_{thr}$ and $\beta$. Just like SNN_3D, the first two layer uses reset-to-zero while the last layer does not reset its membrane potential. In this model, $T_{W}=T_{GT}=St=4$ ms and hence does not require any additional pre-processing as seen in Section III-C; hence, it is called a streaming mode since inputs can stream in directly and continuously to this model.
III-C Input Spike Processing
Figure 5: Input data pre-processing methods presented in this paper: a) Summation mode, where the number of spikes detected within a bin window $T_{W}$ for each probe is summed to create a feature. b) Sub-window mode, where the bin window is further divided into $m$ sub-windows, and the number of spikes detected within each sub-window is summed. c) Streaming mode, where the input spike is gathered as it is.
The spikes generated by the NHP’s neurons are sparse in nature. While SNNs are designed with sparsity in mind, standard neural networks are not and hence require a feature extraction step from the raw spike data. Also, from the biological viewpoint, it is generally assumed that short term firing rates are important for motor control. Hence, we calculate firing rates, $r_{i}(t_{k})$ at the sample time $t_{k}$ from the spike waveforms $P_{i}=\sum_{t_{s,i}}\delta(t-t_{s,i})$ on the i-th probe ($1\leq i\leq N_{ch}$) using the following equation:

### v2

# Combining SNNs with Filtering for Efficient Neural Decoding in Implantable Brain-Machine Interfaces

@@ Abstract
While it is important to make implantable brain-machine interfaces (iBMI) wireless to increase patient comfort and safety, the trend of increased channel count in recent neural probes poses a challenge due to the concomitant increase in the data rate. Extracting information from raw data at the source by using edge computing is a promising solution to this problem, with integrated intention decoders providing the best compression ratio. Recent benchmarking efforts have shown recurrent neural networks to be the best solution. Spiking Neural Networks (SNN) emerge as a promising solution for resource efficient neural decoding while Long Short Term Memory (LSTM) networks achieve the best accuracy. In this work, we show that combining traditional signal processing techniques, namely signal filtering, with SNNs improve their decoding performance significantly for regression tasks, closing the gap with LSTMs, at little added cost. Results with different filters are shown with Bessel filters providing best performance. Two block-bidirectional Bessel filters have been used–one for low latency and another for high accuracy. Adding the high accuracy variant of the Bessel filters to the output of ANN, SNN and variants provided statistically significant benefits with maximum gains of $\approx 5\%$ and $8\%$ in $R^{2}$ for two SNN topologies (SNN_Streaming and SNN_3D). Our work presents state of the art results for this dataset and paves the way for decoder-integrated-implants of the future.
@@ 1 Introduction
Implantable Brain-Machine Interfaces (iBMI) (Figure 1(a)) are a promising class of assistive technology that enables the reading of a person’s intent to drive an actuator[1]. It holds promise to enable paralyzed patients to perform activities of daily living with partial or total autonomy[1]. While the first applications were in motor prostheses to control a cursor on a computer screen[2], or wheelchairs[3], or robotic arms[4], recent studies have shown remarkable results for speech decoding[5, 6], handwritten text generation[7] and therapies for other mental disorders[8].
@@
The majority of clinical iBMI systems have a wired connection from the implant to the outside world [1] that restrict’s the user’s mobility. Many studies have found user independence to be a top priority for patients[9]. In addition, the wired connection also entails a risk of infection leading to an increasing interest in wireless neural interfaces [10, 1, 11]. Another trend in the field has been the constant increase in the number of electrodes[12] to increase the number of simultaneously recorded neurons (Figure 1(b)) which can increase the precision of decoding the intent of the user and enable dexterous control. The recently developed Neuropixels technology has increased the number of recorded neurons to $\approx 1000$. This can be problematic for wireless implants due to the conflicting requirements of high data rate and low power consumption [13]. Hence, there are efforts to compress the neural data on the sensor by extracting information from it by edge computing (Figure 1(c,d)).
@@
Different degrees of computing can be embedded in the implant, from spike detection, classification, to decoding[14]. When the occurrence of spikes is only of interest, spike detection methods can be used to only transmit the occurrence of spikes and can provide compression rations between 100-1000x (depending on firing rate and signal to noise ratio) compared to the conventional data rate [14, 15]. However, this method removes all waveform information that could enable the detection of source neurons necessary in neuroscientific experiments. Spike sorting is another compression method in which in addition to the spike, an identifier of the firing neurons is also sent, resulting in slightly reduced compression rates[14, 16]. Ideally, decoding on implant can provide the maximum compression [17] with the added benefit of patient privacy since the data does not need to leave the implant—only motor commands are sent out. Figure 1(c) compares the transmission data rate of three methods [18, 17] assuming 100/1000/10000 channels, neural firing rate of 100 Hz, two classes and decoder output rate of 250 Hz. As the number of channels increases, decoding offers the best compression, as its output data rate is fixed (albeit at the cost of increased decoder complexity). Traditional decoders have used methods from statistical signal processing such as Kalman filters and their variants[19, 20]. With the rapid growth of Artificial Neural Networks (ANN) and variants for many different applications due to their natural ability to model nonlinear functions and availability of special hardware for training, it is natural to explore the usage of such techniques for motor decoding and several such works have recently been published [17, 21, 22, 23, 24].
Figure 1: (a) Conceptual figure of an iBMI that reads user intent and controls an effector. (b) Trend of exponential increase in the number of simultaneously recorded neurons[25]. (c) Transmission data rates for the different cases (Raw data, spike sorting, decoding). While spike sorting provides some compression, decoding on the implant can provide the best option, especially as the number of electrodes increases beyond 1000. (d) Integrating computing in the implant can reduce the wireless datarate enabling scalability of iBMI systems.
@@
To fit on the implant, the decoder has to be extremely energy and area efficient, along with being accurate. Some specialized decoder integrated circuit using ANNs have been developed to achieve this purpose[18, 22, 24]. Brain-inspired SNN are supposed to be even more energy-efficient due to their event-driven nature [26, 27, 28]. They are also expected to be better at modeling signals with temporal dynamics due to their inherent “stateful” neurons with memory. However, detailed comparisons between SNN and ANN variants with controlled datasets and benchmarking procedures have been lacking. A recent effort[29] has put together a benchmarking suite to address this gap and one chosen task is that of motor decoding. We use the same dataset for benchmarking and show additional results for more control cases.
@@
Neurobench showed that streaming SNNs provide a good tradeoff in terms of accuracy vs computes while other methods could have similar memory footprint. Another recent work[30] using the same benchmarking suite showed recurrent networks providing the best results with LSTMs outperforming SNNs in terms of accuracy. We make the following novel contributions in this paper:
  • We improve SNNs for regression by filtering the output using a Bessel filter from signal processing and close the gap in accuracy with LSTMs.
  • We demonstrate that filtering the outputs of both ANN and SNN decoders with block Bidirectional Bessel filters improves decoding accuracy.
  • We demonstrate state-of-the-art decoding accuracy on this benchmark using LSTMs and filtered SNNs which occupy the higher and lower ends of the pareto optimal curve respectively.
@@
  • We show improved accuracy when trained with better curated data; this points to an automated method of trial selection for training neural networks for neural decoding.
The rest of the paper is organized as follows. The following section discusses some of the related works while Section 3 describes the dataset, models and pre-processing used in this work. Section 4 presents the results comparing different models in terms of their performance-cost tradeoff using pareto curves. This is followed by a Section 5 that discusses the main results and provides additional control experiments. Finally, we summarize our findings and conclude in the last section.
2 Related Works and Contribution
@@
2.1 Traditional Signal Processing Decoders
@@
2.2 Machine Learning Decoders: Algorithms
@@
Neuromorphic algorithms have emerged as an energy-efficient decoder and an effective tool for data compression [39]. SNN is a brain-inspired neural network popular in neuromorphic applications due to its low energy. It can achieve nearly the same accuracy as ANN but with less than $10\%$ memory access and computation of ANN [40]. Similarly, it was found that the SNN decoders could use far fewer computes compared to ANN, but with a performance penalty in accuracy, for the motor prediction of primates in the Neurobench benchmark suite[29]. Another recent work [30] showed that SNNs offer an advantage of low-latency which is essential for closed-loop neuromodulation.
In this work, we show that the combination of traditional signal processing filters with SNNs results in one of the best decoders for iBMI systems.
2.3 Machine Learning Decoders: Hardware
Specialized hardware implementations are needed to fit machine learning decoders within the strict area and power budgets of implants. One of the earliest works[18] used a hardware-algorithm co-design approach to exploit statistical variations in analog circuits to make a sub-microwatt decoder based on extreme learning machine, a variant of reservoir computing algorithms. It also used a configurable digital processing second stage to program distinct weights learned for each chip. More recent work[24] has used general purpose M0 processor and digital matrix acceleration units, where the power efficiency stems from usage of special features called spike band power (SBP) that require much lower sampling rate than conventional spike detection. While these earlier works demonstrated promising decoder hardware, they were not integrated with neural recording amplifiers as a system on chip. This has been a recent focus[22] where multiplexed neural recording front-end circuits are integrated with a 31-class decoder for brain to text applications. In this case, the classifier was as simple linear discriminant analysis (LDA) but its performance is enhanced by a preceding feature extraction module that extracts distinct neural codes. While all of these works used traditional ANNs, one example[27] used a SNN decoder to perform a closed loop decoding task of moving an object to a desired location using rodents. Based on the decoded outputs in each step, intra-cortical micro stimulation was delivered to close the loop and allow error correction over multiple time steps. This was a multi-component system not optimized for power dissipation. There is significant opportunity to improve SNN hardware and integrate with neural amplifiers to create a system on chip.
3 Methodology
@@
  • $T_{GT}$: The fixed time interval for ground truth labels.
@@
3.1 Dataset
The primate reaching dataset chosen for this paper was gathered and released by [41], with the six files chosen for Neurobench [29] being the files of interest. These six files are recordings of two non-human primates (NHP) (Indy and Loco), where each NHP accounts for three files (more details about this choice in [29]).
This dataset contains microelectrode array (MEA) recordings of the NHP’s brain activity while it is moving a cursor to the target location, as seen in Figure 2. The finger velocity is sampled at $f_{GT}=250$ Hz resulting in ground truth labels at a fixed interval of $T_{GT}=4$ ms. The target position changes once the monkey successfully moves the cursor to the intended target. We refer to this action as a reach. The dataset contains a continuous stream of the brain’s activity from one MEA with $N_{ch}=96$ probes (Indy) or two MEAs with $N_{ch}=192$ probes (Loco). In this work, we ignore sorted spikes since it has been shown that spike detection provides sufficient information for decoding[18, 42] and is more stable over time. Hence, the number of probes $N_{ch}$ is the input feature dimension $N_{0}$ for the neural network models (except ANN_3D) that will be discussed in the following subsection.
@@ 3.1 Dataset
Training NN models on time series-based data requires the data to be split apart into separate segments. In analogy with keyword spotting[29], each segment of neural data should correspond to separate keywords. By using the target positions in this dataset, we can separate the spike data into segments based on indices in the target position array where there is a change in values, as illustrated in Figure 3. Such consecutive indices forms the beginning and end of a reach, and then we can split the time series into training, validation, and test sets based on the number of reaches. The split ratio used in this paper follows that of Neurobench[29], which is 50% for the training set, and 25% each for validation and test sets. The total number of reaches recorded in each file can be seen in Table 1.
@@
Table 1: Number of reaches in each file [71]
@@
indy_20160622_01 970
indy_20160630_01 1023
indy_20170131_02 635
loco_20170210_02 587
loco_20170215_02 409
loco_20170301_05 472
3.2 Network Models
To explore the potential of various neural network models as the neural decoder, five different model architectures with and without memory are tested: ANN, ANN_3D, SNN_3D, Streaming SNN and LSTM, which can be seen in Figure 4. These five models use NN architectures popular as neural decoders (e.g. ANNs used in [43] [44], SNNs used in [26] [45] [46] [47] and LSTMs used in [48] [49] [50] [51]) and have memory at the input layer or hidden layer. Every model except for LSTM has two versions of varying complexity (explained in Section 4.1) where complexity refers to the model size indicating the number of neurons. The larger model is henceforth referred to as the base model while the smaller model is dubbed the tiny variant. It was found that networks deeper than 3 layers performed poorly and hence deeper models were excluded from this study.
@@ 3.2 Network Models
    The ANN model has an architecture of $N_{ch}-N_{1}-N_{2}-2$, with rectified linear unit (ReLU) as the activation function for the first two-layers as well as batch normalization to improve upon the accuracy obtained by the model. Note that $N_{0}=N_{ch}$ indicates one feature extracted from each probe obtained by summing the neural spikes over a fixed duration of $T_{W}$ as described in Section 3.3. Also, $N_{3}=2$ corresponds to predicting the $X$ and $Y$ velocities. A dropout layer with a dropout rate of $0.5$ is also added to the first two layers to help regularize the model. In analogy with the naming convention of ANN_3D introduced next, this model can also be referred to as ANN_2D due to the shape of the input weight tensor.
@@
    The architecture of the ANN_3D or ANN_flat model is $m\times N_{ch}-N_{1}-N_{2}-2$, i.e. it shares an identical architecture with ANN, except at the input layer. This model divides the $T_{W}$ duration of the input bin window into $m$ sub-windows and creates a $m$-dimensional feature from each probe by summing spikes in each sub-window. This mode of input will be further explained in Section 3.3. The input will then be flattened across the sub-windows, yielding a final input dimension of $N_{ch}\times m$; hence, the number of weights/synapses in the first layer is $m$ times more than ANN. It is referred to as ANN_flat in [29]; we refer to it as ANN_3D here in keeping with the shape of the input weight tensor, which we feel is more intuitive.
@@
    The SNN_3D aims to achieve high accuracy and shares a similar architecture with ANN ($N_{ch}-N_{1}-N_{2}-2$), with the following differences: 1) Instead of using standard activation like ReLU, the SNN_3D model uses the leaky integrate-and-fire (LIF) neuron after every fully-connected layer, 2) the input is first passed through layer normalization, similar to LSTM due to the recurrent nature of LIF, 3) at the final layer there is a scaling layer applied to the output LIF neurons and 4) the input spikes are processed using the sub-window method similar to ANN_3D to capture finer temporal details of the spike rate variations. However, the dimension of the input layer is not $m\times N_{ch}$ like ANN_3D since in this case, the spikes from the $m$ sub-windows are fed over $m$ time steps to $N_{ch}$ neurons in the first layer using a single weight synapse. The LIF neurons are governed by the following set of equations:
@@
    where $U[t]$ and $X[t]$ are the membrane potential of the LIF neuron and the input at the t-th time step respectively, $W$ is the synaptic weight of the fully-connected layer, $\beta$ is the decay rate, $S_{out}[t]$ is the output spike, $U_{thr}$ is the membrane potential threshold, $U_{sub}$ is the subtracted value if the reset mechanism is reset-by-subtraction and $\theta$ is the reset mechanism. The LIF neurons for all layers shares the same $U_{thr}$ and $\beta$. The first two layer uses the reset-to-zero mechanism while the last layer does not use any reset to allow the final output neurons to accumulate membrane potential to predict the velocity of the primate’s movement. For every stride of $4$ ms, the membrane voltages are reset and the integration is restarted with fresh input to produce the next output. Due to the reset of the LIF neurons after every prediction, overlapping bin-windows (for $T_{W}>St$) cause the SNN_3D to process the same input spikes for multiple predictions.
@@
    The SNN_Streaming model also consists of three fully-connected layers ($N_{ch}-N_{1}-N_{2}-2$), with LIF neurons (Equation 1) in each layer. Unlike SNN_3D, every LIF layer has its own unique $U_{thr}$ and $\beta$. SNN_3D is designed to achieve the highest accuracy while SNN_Streaming is designed to achieve the best tradeoff between accuracy and resource consumption. Accordingly, the two main differences between SNN_3D and SNN_streaming are the usage of Layer Normalization and the input data processing. SNN_Streaming avoids using Layer Normalization, which removes data sparsity (by subtracting the mean from the 0 values). The resulting model of SNN_Streaming has higher sparsity and enables the construction of an energy-efficient model by reducing computations. In terms of input spike processing, $T_{W}=T_{GT}=St=4$ ms in this model and hence it does not require any additional pre-processing as seen in Section 3.3; hence, it is called a streaming mode since inputs can stream in directly and continuously to this model. Just like SNN_3D, the first two layer uses reset-to-zero while the last layer does not reset its membrane potential.
3.3 Feature Extraction by Input Spike Processing
Figure 5: Input data pre-processing methods for feature extraction presented in this paper: a) Summation mode, where the number of spikes detected within a bin window $T_{W}$ for each probe is summed to create a feature. b) Sub-window mode, where the bin window is further divided into $m$ sub-windows, and the number of spikes detected within each sub-window is summed. c) Streaming mode, where the input spike is gathered as it is.
The spikes generated by the NHP’s neurons are sparse in nature. SNNs can intrinsically accept sparse spiking input since they create an accumulation in the membrane potential variable. For ANNs however, the information over a past time period has to be explicitly accumulated in a feature extraction step. Also, from the biological viewpoint, it is generally assumed that short term firing rates (as opposed to mean firing rates over a trial duration [52]) are important for motor control [53] [18]. Hence, we calculate firing rates, $r_{i}(t_{k})$ at the sample time $t_{k}$ from the spike waveforms $P_{i}=\sum_{t_{s,i}}\delta(t-t_{s,i})$ on the i-th probe ($1\leq i\leq N_{ch}$) using the following equation: